OpenStudy (anonymous):

Which fraction is between 1/2 and 1? A.1/11 B.2/12 C.8/9 D.1/5

OpenStudy (anonymous):

help

OpenStudy (anonymous):

lets try, 1/2= 0.5

OpenStudy (anonymous):

1/11= 0.09 2/12=0.16 8/9=0.88 1/5=0.2

OpenStudy (anonymous):

The answer would be C because of what we stated up above